OpenStudy (anonymous):

Can some one help me with just this step: Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4).

OpenStudy (anonymous):

look at my first question for volume, length, width, height

hero (hero):

If it was morning instead of night, I'd probably be in a better helping mood.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I just need help in writing the formula..

OpenStudy (anonymous):

Ok I'm looking at it now. First you need to round to integer values. (see line 2) length = x = 8 width = 1 height = 11 I think you need to leave it as V=(1)(11)x Then replace the (1) with (x-7) and the (11) with (x+3) so it will be V=(x-7)(x+3)x. What do you think?

OpenStudy (anonymous):

Horrible question, by the way - hard to know where they are going with it.

OpenStudy (anonymous):

yeah..rounding is a good idea..

OpenStudy (anonymous):

yeah its a very bad question

OpenStudy (anonymous):

The fourth paragraph is odd. Do they mean 77=(x-7)(x+3)x ?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

then maybe you expand and put it as 0=x^3 etc

OpenStudy (anonymous):

oh ok..I got it now..thanks for helping...

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

lol